An open container holds ice of mass 0.525 kg at a temperature of -10.6 ∘C . The mass of the container can be ignored. Heat is supplied to the container at the constant rate of 760 J/minute .

The specific heat of ice to is 2100 J/kg⋅K and the heat of fusion for ice is 334×103J/kg.

Part A How much time tmelts passes before the ice starts to melt?

tmelts = ______________ minutes

Part B From the time when the heating begins, how much time trise does it take before the temperature begins to rise above 0∘C?

trise = ________________ minutes
Answer #1

a) P in = m c dT

760*t = 0.525*2100*10.6

t = 15.377 min

b)now P in = m c dT + m L

760*t = 0.525*2100*10.6 + 0.525*334.0E3

t = 246.1 minutes